OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4). 6.

OpenStudy (anonymous):

First you must find the slope. Do you know how?

OpenStudy (anonymous):

y2-y1 divided by x2-x1

OpenStudy (anonymous):

y2-y1 ----- x2-x2 4-4 ---- 5-2 = 0

OpenStudy (anonymous):

then you do y=mx+b

OpenStudy (anonymous):

plug in one of the coordinates ill use (2,4)

OpenStudy (anonymous):

4=0(2)+b

OpenStudy (anonymous):

4=b

OpenStudy (anonymous):

so 4=b is my answer?

OpenStudy (anonymous):

you have to put it in y=mx+b form

OpenStudy (anonymous):

since x is 0 that mens there is no x so it will be y=4

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

u think u can help me solve another